OpenStudy (anonymous):

Please help. What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10 (−4, −2) and (2, 5) (−2, −4) and (2, 5) (−2, −4) and (5, 2) No Real Solutions

OpenStudy (anonymous):

the second equation is the same as \[y=8x-10\] plug it in to the top one and you can solve \[8x-10=x^2+12x+30\]

OpenStudy (anonymous):

I got \[x ^{2}+4x+40\]

OpenStudy (agl202):

No Real Solution!!!!!

OpenStudy (anonymous):

set it equal to zero via \[x^2+4x+40=0\]

OpenStudy (anonymous):

yeah like that

OpenStudy (solomonzelman):

subtract 36 from both sides, and complete the square.

OpenStudy (anonymous):

and then @Agl202 tell you that there is no real solution, which is true

OpenStudy (agl202):

sorry

OpenStudy (solomonzelman):

you were right, alg.

OpenStudy (agl202):

thx....

OpenStudy (solomonzelman):

I was going to say, that from where satelline left off, if you subtract 36 from both sides, and complete the square (as i said before), you will then end up with imaginary solutions.